Find the perimeter of a rectangle, given that it's length is 7 and the length of it's diagonal is 25. Round your answer to the nearest hundredth
Answer by stanbon(75887)

Find the perimeter of a rectangle, given that it's length is 7 and the length of it's diagonal is 25. Round your answer to the nearest hundredth
------------------------
Let the width be "w"
---
Equation:
w^2 + 7^2 = 25^2
w^2 = 625-144
w^2 = 481
-------------------
w = 21.93
---------------------------------
Perimeter = 2(length + width)
P = 2(7 + 21.93)
P = 2(28.93)
P = 57.86
====================
Cheers,
Stan H.